A cylinder of diameter 1cm at temp 30c is to be slid into a hole in steel plate.the hole had a diameter of 0.99970cm at temp 30c.To what temp must the plate be heated?


1
Expert's answer
2021-03-14T19:21:51-0400

We can find to what temperature must the plate be heated from the thermal linear expansion formula:


"\dfrac{\Delta L}{L_0}=\alpha \Delta T,""\Delta T=\dfrac{L-L_0}{\alpha L_0},""\Delta T=\dfrac{1\ cm-0.9997\ cm}{12\cdot10^{-6}\ ^{\circ}C^{-1}\cdot0.9997\ cm}=25^{\circ}."